OpenStudy (anonymous):

Which of the folowing is both ariphmetic and geometric series? 1) 0 1 2 3 4 5 6 2) 1 2 4 8 16 32 3) 2 2 2 2 2 2 2 4) 1 1 2 3 5 8 13

Nnesha (nnesha):

ariphmetic ???

OpenStudy (anonymous):

there is no such series, one that is both geometric and arithmetic

OpenStudy (solomonzelman):

Difference zero (arithmetic) Common ratio 1 (geometric). For example option 3. LOL

OpenStudy (anonymous):

what is geometricc and ariphmetic serieses? I don't know

Nnesha (nnesha):

that suppose to be arithmetic so geometric when terms are increasing or dividing by same number arithmetic when terms are adding or subtracting by same number

Nnesha (nnesha):

multiplying**** not increasing

OpenStudy (anonymous):

i gues the 2 2 2 2 2 2 2 2

OpenStudy (anonymous):

ok, so the answer is 2 2 2 2 2 2 2. Thank you

OpenStudy (anonymous):

i am fairly sure that you do not consider the constant series geometric, but i guess i could be wrong

OpenStudy (solomonzelman):

Well, I wouldn't think there is REALLY such a thing as a series that is arithmetic and geometric at the same time though. Aren't the numbers supposed to be changing somehow ?

OpenStudy (anonymous):

@EmmaMink

OpenStudy (anonymous):

sort of like saying \[f(x)=1^x\] is exponential

OpenStudy (solomonzelman):

LOL, yeah:) (that can be a proof that all numbers are equal. 1^3=1^4=1^0 ... 3=4=0 )

OpenStudy (emmamink):

And yes, 2222222 is correct. Arithmetic and Geometric sequences are true when you add the same number and subtract the same number. Check: 2 + 2 - 2 = 2!

OpenStudy (solomonzelman):

So for 0 0 0 0 Check: 0 + 0 - 0 = 0! ?

OpenStudy (emmamink):

I gave you a medal ttb123456789

OpenStudy (anonymous):

another $20 says this was written by the morons at FLVS

OpenStudy (solomonzelman):

Emma your conclusion gives that 2 2 2 2 2 is arithmetic and geometric, but 0 0 0 0 0 is what then ?

OpenStudy (emmamink):

http://www.purplemath.com/modules/series3.htm

OpenStudy (emmamink):

I found a site that explains this.

OpenStudy (anonymous):

i have no idea why the state of florida condones this kind of nonsense i guess they really do not know any better

OpenStudy (anonymous):

i;m in WI

OpenStudy (emmamink):

Well whatever party people Emma is out! Peace!!!

OpenStudy (anonymous):

I think no sites can be better than good math solvers. you are all good math solvers

OpenStudy (emmamink):

Well if I'm good give me medal. EMMA WANT MEDAL!!!!! GIMME!!!

OpenStudy (emmamink):

jk i dont deserve 1

OpenStudy (solomonzelman):

@EmmaMink ASKING FOR MEDALS DIRECTLY, IS A VIOLATION OF THE oPENsTUDY POLICY!

OpenStudy (solomonzelman):

unless it is a joke of course. But then how do you know ?

OpenStudy (anonymous):

1+1+1+1+1+1+.........is both arithmetic and geometric

OpenStudy (emmamink):

Cause I said jk I don't deserve one? jk means just kidding.

OpenStudy (solomonzelman):

Surjithayer, how about 0 + 0 + 0 + 0 + 0 ?

ganeshie8 (ganeshie8):

i think 0 is not allowed as a term in geometric series

OpenStudy (anonymous):

how to calculate common ratio?

ganeshie8 (ganeshie8):

so your question makes no sense

OpenStudy (solomonzelman):

Yes, ganeshie that would make sense, because multiplying zero wouldn't change it.

OpenStudy (solomonzelman):

Wait, how about negative integers?

ganeshie8 (ganeshie8):

see the definition http://mathworld.wolfram.com/GeometricSeries.html

OpenStudy (anonymous):

1+1+1+1+1+1+.........is both arithmetic and geometric i would disagree first of all, it is not a sequence it is a series

ganeshie8 (ganeshie8):

if the ratio of two consecutive terms is constant then it is geoemtric. the definition is not particular about negatives/postives... but you cannot have 0 as a term because the geometric series is based on "ratio" of consecutive terms

OpenStudy (anonymous):

secondly if you mean the series of partial sums, they are \[1,2,3,4,5,...\] which is a sequence, but not a geometric one

OpenStudy (anonymous):

oops make that "sequence of partial sums"

ganeshie8 (ganeshie8):

1+1+1+1+1+1+... meets the definition of geometric series right ?

OpenStudy (anonymous):

\[1+ 1 + 1 + 1 + 1...\] has no meaning i can discern \[\sum_{k=1}^{\infty}k\] is not a number

ganeshie8 (ganeshie8):

maybe lets call it degenerated series or whatever, but i don't see how it violates the definition